If you receive a negative value for a variance and standard, you must check your computations. Is this statement true or false?
1
Expert's answer
2020-08-17T18:05:18-0400

This statement is true because a variance (and a standard deviation) cannot be negative.

Let we have a r. v. "\xi". Then "D\xi=M(\xi-M\xi)^2". So we consider a r. v. "(\xi-M\xi)^2" (which takes on only non-negative values) and find its expectation. We get "D\xi\geq 0".

"D\xi=0 \Leftrightarrow \xi-M\xi=0\ (\text{when }\xi \text{ is constant})."

"\sigma_\xi=\sqrt{D\xi}". So "\sigma_\xi\geq 0".